OpenStudy (anonymous):

solve: (6x-4)^1-4 = (5+3x)^ 1/4

OpenStudy (anonymous):

(6*x-4)^1-4 = ((3*x+5)^1)/4 // - ((3*x+5)^1)/4 (6*x-4)^1-(((3*x+5)^1)/4)-4 = 0 6*x-((3*x+5)/4)-4-4 = 0 (-1*(3*x+5))/4+6*x-4-4 = 0 (-1*(3*x+5))/4+(4*6*x)/4+(-4*4)/4+(-4*4)/4 = 0 4*6*x-1*(3*x+5)-4*4-4*4 = 0 21*x-16-16-5 = 0 21*x-21-16 = 0 21*x-37 = 0 (21*x-37)/4 = 0 (21*x-37)/4 = 0 // * 4 21*x-37 = 0 21*x-37 = 0 // + 37 21*x = 37 // : 21 x = 37/21 x = 37/21